Mal87:

If a gas in a closed container is pressurized from 15.0 atm to 16.0 atm and its original temperature was 25 °C, what would be the final temperature of the gas in Kelvin? A. 279 K B. 299 K C. 318 K D. 338 K

Vocaloid:

late reply but pressure is proportional to temperature P1/T1 = P2/T2; note this only works if all temperatures are converted to Kelvin first so 15/(25+273) = 16/T2 should give you your second temp